Identified OR, RR and the differences between them. Can you please elaborate on their differences in terms of type I error?
Answer #1

ODD RATIO : This can be defined as a measure of association between an exposure and an outcome. Here the odd ratio represents the odds that an outcome will occur with a given particular exposure

or

This can be defined as the ratio of the odds of an event occuring in one group to the odds of occuring in another group

Relative Risk : This can be defined as the risk of an event relative to the exposure

The main basic difference is that the odds ratio is a ratio of two odds and   whereas the relative risk is a ratio of two probabilities.  

Formula For Odd Ratio :

Odd of event in Treatment Group / Odds of event in control group

= a / b / c / d

= ad / bc

Formula For Relative Risk :

Risk of the event in treatment group / Risk of the event in the control group

= a / (a + b ) / c / (c + d)